# Improving Multimodal Large Language Models Using Continual Learning

### v1

@@ 3.2 Mitigating Linguistic Forgetting
We study the efficacy of our mitigation methods toward reducing linguistic forgetting. Due to computational constraints, we exhaustively tested mitigation methods with Pythia (160M) in Table 3(a), and then evaluate the best method across all parameter scales of Pythia models in Fig. 3(b).
In Table 3(a), Soft Targets has the highest accuracy across VL datasets with the least linguistic forgetting. LoRA and mSGM better preserve NLU/NLG performance but at the cost of decreased VL accuracy compared to naive fine-tuning. We posit that the retention of next-token logits enforced by Soft Targets is crucial in mitigating linguistic forgetting for causal generation during MLLM training.
@@
VQAv2 TextVQA GQA Acc $\uparrow$ $\Delta\downarrow$ Acc $\uparrow$
OCR Pure
@@
(a) Results for mitigation methods on LLaVA training. Evaluate alternatives to naive fine-tuning for transforming Pythia 160M into an MLLM. In Fig. 3(b) (right), we evaluate the average linguistic forgetting for Pythia models in the 160M to 2.8B scale, after training on LLaVA. Negative forgetting refers to a positive backward transfer in NL performance after multimodal training.
(b) Linguistic forgetting for varying model sizes of Pythia (160M - 2.8B).
Analyzing the Role of Parameter Count. In Driess et al. [6], larger models had less catastrophic forgetting than smaller ones for NLG/NLU. The opposite result was found in Luo et al. [30].
We analyze this phenomenon in the Pythia family of models in Fig. 3(b), where we evaluate naive fine-tuning and the best method for 160M, Soft Targets. Across scales, Soft Targets has zero or negative linguistic forgetting over the naive fine-tuning method used in the original LLaVA paper, with competitive VL performance compared to naive fine-tuning (especially at higher scales, see Fig.  4(a)). Soft Targets achieved positive backward transfer in the low model size regime (0.16-0.41B) and no forgetting for larger sizes (>0.41B). In contrast, naive fine-tuning had severe forgetting in the 0.16-0.41B regime and reduced forgetting when the model size exceeded 1B parameters.
@@ 3.3 Continually Learning VL Tasks
We next turn to continually learning each of the VL datasets used to train LLaVA 1.5, where we group the datasets based on the task type. Details of the data mixture, task groupings and ordering are provided in Table 4 and Sec. B.3 in the Appendix. Note that there is no VL evaluation dataset associated with the Task 2 (Instruct), but we still measure NLU/NLG performance. Given our limited computational budget, we exhaustively studied our CL methods only for the smaller-scale Pythia 410M LLM (see Table 1), and then we evaluated the best-performing mitigation method for all of the Pythia LLMs (see Appendix Figures  5(a) and  5(b)).
Table 1: Continually learning LLaVA Tasks with Pythia 410M. We report cumulative task-wise accuracy and forgetting of each mitigation method across VL and NL tasks, where we evaluate test sets associated with all tasks seen up to current task. Task 5 represents cumulative performance.
@@
Rehearsal $(1\%)$ 0.58 37.74 10.90 3.47 7.41 3.55 7.65
mSGM + Reh. $(1\%)$ 1.11 35.28 0.73 12.38 2.25 10.21 2.77
@@
Results for Pythia 410M are given in Table 1. In terms of NLU/NLG forgetting, all mitigation methods showed efficacy in reducing forgetting across the sequence of tasks. Overall, mSGM with and without rehearsal achieves the least linguistic forgetting while maintaining the highest cumulative VL performance across all baselines. Regarding VL performance, naive fine-tuning achieves better performance at the cost of high linguistic forgetting. In contrast, mSGM achieves competitive VL performance and sometimes even surpasses it, e.g., for RefCOCO, an especially challenging VL task.
Model scaling results. Model scaling results are given in Appendix Fig. 5. We compare mSGM with Rehearsal - the best method identified for Pythia 410M, against naive LLaVA fine-tuning across all Pythia model sizes. In general, the original naive LLaVA fine-tuning exhibits greater loss of NLU/NLG as well as VL performance. In contrast, VL performance for mSGM rivals or exceeds the original LLaVA fine-tuning across all model sizes, with little to no NLU/NLG forgetting.
@@ 4 Discussion
This work presents one of the first studies of linguistic forgetting in MLLMs, particularly for open-source models with modest parameter counts ($<7$B). We show that the degree of linguistic forgetting typically reduces with model scale, and is far more severe for NLG tasks compared to NLU. We pioneer treating MLLM creation as a CL problem, and show that CL methods effectively mitigate linguistic forgetting while minimally hindering VL accuracy. In our experiments, our best mitigation method far outperforms the naive LLaVA recipe in terms of linguistic forgetting, while maintaining competitive VL performance. We show that this benefit holds across model scales. We pioneer CL for MLLMs, and establish strong baselines for this task. In fact, we observe that besides maintaining competitive VL performance with naive LLaVA training, our mitigation methods achieve near zero and in some cases even negative linguistic forgetting. This suggest that our mitigation methods achieve a positive backward transfer of linguistic ability after multi-modal training. Given the essential nature of multimodal abilities for many applications, our research highlights that naive multimodal fine-tuning can significantly degrade prior linguistic abilities. Our findings underscore the need to develop more robust and capable mitigation approaches, and showcase the applicability of CL techniques to the fine-tuning of foundation models. We aim to inspire future research in this direction, ultimately contributing to the advancement of more resilient and versatile MLLMs.
@@ Acknowledgments
  • Appendix A describes the implementation details for the experiments.
  • Appendix B provides details on the datasets and evaluations.
@@
  • Appendix C summarizes findings of additional supporting experiments and ablation studies.
Appendix A Implementation Details
We provide details below of the experimental settings for the alignment and fine-tuning phases of the LLaVA1.5 training protocol, the continual learning methods methods we employ, and the implementational and hyperparameter search details. For LLaVA, our configuration details are the same as LLaVA1.5 in order to reproduce and compare effectively to the LLaVA1.5 multi-modal training protocol.
A.1 Libraries and Tools
To reproduce the LLaVA 1.5 experiment, we build on top of the Prismatic library [31], and use the Prismatic library for vision-language tasks and EleutherAI’s LM_Eval library [32] for natural language evaluations. All our code is written in PyTorch [33].
A.2 Base LLMs used
We study 9 choices for LLaVA’s LLM: comprising six Pythia models [17], Phi2 (3B)[18], and two LLaMA 2 (7B) models. The six Pythia variants span various scales, including 160M, 410M, 1B, 1.4B, and 2.8B parameters, with two versions of the 1.4B model—one of which is instruction fine-tuned. Additionally, we evaluate both the original LLaMA 2 and the instruction fine-tuned Vicuna-1.5 7B[5], which is utilized in LLaVA 1.5.
  1. 1.
    Phi2 [34] has been trained on the same dataset as Phi1 [18], which includes a curated selection of “textbook quality” data from the web (6 billion tokens) and an additional 1 billion tokens of synthetically generated textbooks and exercises created using GPT-3.5 [35]. The Phi series are among the most performative models in the under 7B parameter size class [34].
  2. 2.
    Pythia [17] comprises two sets of 8 models, each corresponding to two datasets. For every model size, one set is trained on the Pile dataset [36], while the other set is trained on a version of the Pile where global de-duplication has been applied. The granular model scaling suite of Pythia is particularly useful for our study. We select the de-duplicated set of Pythia models.
  3. 3.
    LLaMA 2 is reportedly trained on a mix of publicly available online data, but specific details are not available [37]. LLaVA 1.5 reports its best performance with instruction fined-tuned LLaMA 2 LLMs [12].
A.3 LLaVA Training Details
Following the LLaVA 1.5 protocol  [8], the visual encoder is a CLIP VIT-L@336px which takes an image resolution of 336px with letterbox resizing, and the alignment network is a two-hidden layer MLP projector with GELU activation. The LLM generations are limited to a maximum token length of $2048$.
A.3.1 Alignment Stage
In the alignment stage, we use a learning rate of $0.001$ with a linear warmup followed by a cosine decay scheduler. The training process is carried out for $1$ epoch with a global batch size of $256$, distributed as $16$ samples per device. We use gradient checkpointing and mixed precision training to speed up computation. We used FSDP (Fully Sharded Data Parallel) to train all our models [38]. During the alignment stage, we shard only the gradients and optimizer states, not the parameters.
A.3.2 Fine-tuning Stage
For the fine-tuning stage, the learning rate is $052e-$ 05, with a linear warmup for a $0.03$ ratio of steps followed by cosine decay. This stage also runs for 1 epoch but with a smaller global batch size of $128$, keeping the per-device batch size constant at $16$. Depending on the size of the LLM (Phi2 3B vs Pythia 160M), we vary the per-device batch size, but the global batch size is kept constant across all experiments. We switch to the FSDP full sharding strategy, with all parameters, optimizer states, and gradients sharded across the devices.
A.4 Continual LLaVA
For the Continual LLaVA setting, we follow the same experimental configuration as before in the LLaVA setting. A model trained on Task $(i)$, will then be simply used as the pretrained checkpoint for Task $(i+1)$, with training following normally. The training configurations are then identical to the fine-tuning stage. In the case of PET methods like LoRA, the adapter weights are merged back into the LLM before starting training on the new task.
Different from the LLaVA setting, we continually learn each of the VL datasets used to train LLaVA 1.5, where we group the datasets based on the task type. Details of the dataset sequence are described in Appendix 4.
A.5 Continual Learning Methods
To mitigate catastrophic forgetting in MLLMs, we examine and adapt several methods for LLaVA multimodal training:
  1. 1.
  2. 2.
    LoRA keeps the original LLM weights frozen and learns low-rank updates for them [20]. Following LoRA’s recommended protocol, we inject LoRA weights into all LLM linear layers. After each task is learned, LoRA weights are merged into the LLM. Details in Appendix A.6.2.
  3. 3.
    Soft Targets was proposed in [21] to reduce forgetting in CL. Rather than using hard targets for training, simple label smoothing is used to reduce the severity of the training loss under distribution shift in CL. We smooth the hard target vector $Y$, by smoothing the target tokens by $-\alpha$, and offsetting non-target tokens by $+\alpha/(N-1)$, where $\alpha$ controls the smoothing, and $N$ is the LLM’s vocabulary size. We discuss the choice of $\alpha$ in Appendix A.6.3.
  4. 4.
    Rehearsal (Experience Replay) is an effective method for CL that involves storing data from earlier tasks and mixing it with data from new tasks. We study it in our CL experiments since we do not have any stored data for task 1 (training the base LLM). We study storing 1% of randomly selected samples from each previous task, excluding task 1.
  5. 5.
    mSGM is based on SGM, which combines soft targets, weight initialization, and LoRA to mitigate catastrophic forgetting [21]. We adapt SGM based on our modified soft targets and omit weight initialization and output layer freezing since the output vocabulary is static, and the output layers are used for the causal generation of LLMs.
A.6 Hyperparameter Search
A.6.1 LLaVA1.5 Setting
@@ A.6 Hyperparameter Search
To reproduce and directly compare against the LLaVA 1.5 protocol, we keep the explicit training configurations the same (as mentioned in Appendix A).
A.6.2 LoRA
@@
We train several different LoRA settings on the smallest Pythia-160M LLM for tractability and compare the resulting VL and NLU/NLG performances. We vary the a) target modules (between 1) all linear layers, and 2) key, query, and value projection layers only), b) LoRA ranks in the range of $1/4-1/2$ of the full rank of the model, rank stabilized LoRA [39], and larger alpha values ($16$ instead of $8$ as default). Table 2 shows the comparisons by training on the LLaVA setting.
@@
A.6.3 Soft Targets
We train the LLaVA recipe with the soft targets with varying alpha $\alpha\in\{0.001,0.01,0.1\}$, and report results in Figure  3. The Pythia 160M model is used for this tuning.
@@ A.6.3 Soft Targets
Appendix B Datasets & Evaluation
B.1 Training Datasets
For the continual LLaVA setting, the original LLaVA 1.5 data mixure is split into groups of vision-language (VL) tasks. These VL tasks are then learned sequentially in the Continual LLaVA training. Table 4 provides the splits of the LLaVA 1.5 data mixture into a sequence of VL tasks, based on task types. The collation of all these datasets forms the LLaVA 1.5 data mixture, which is used to train the LLaVA MLLMs, per the LLaVA’s protocol as described in Appendix  A.3.
Table 4: Continual LLaVA Setup: The LLaVA 1.5 data mixture is split into groups of vision-language (VL) tasks. VQA (OE & OK) refers to open-ended and outside-knowledge VQA tasks.
Task Type Task Data Size
Pre-Training 1 LLM Pre-Training*
Instruct Tuning 2 CC-LAION-SBU 558K
LLaVA-Inst, ShareGPT 198K
VQA 3 VQA2 83K
(OE & OK) OKVQA 9K
A-OKVQA 66K
GQA 72K
VQA (OCR) 4 OCRVQA 80K
TextCaps 22K
Referential 5 RefCOCO 48K
Grounding VisualGenome 86K
B.2 Evaluation Datasets
Natural Language Evaluation: We use six datasets. For NLG, we use Lambada [23], which is the only NLG dataset that uses accuracy for evaluation. For NLU, we use ARC-Easy [24], ARC-Challenge [24], Winogrande [25], and WSC [25] for NLU.
Vision-Language Evaluation: We use the test sets corresponding to each VL dataset used for training LLaVA: VQAv2 and GQA for general VQA tasks [26, 27], TextVQA OCR (and Pure) for OCR tasks [28], and RefCOCO for referential expression generation tasks [29]. Additional details are given in Appendix B.3.
B.3 Evaluation Dataset Preparation

### v2

@@ 4.2 Mitigating Linguistic Forgetting
Following the results in Sec. 4.1, we study the efficacy of our mitigation methods toward reducing linguistic forgetting. Due to limited computational resources, we only exhaustively tested the mitigation methods with Pythia (160M), which was selected due to its small size. We then evaluate the best method using all Pythia models in Sec. 4.2.
Our results with various mitigation methods on LLaVA training of Pythia (160M) are given in Table 3. Soft Targets has the highest accuracy across VL datasets with the least linguistic forgetting. LoRA and mSGM better preserve NLU/NLG performance but at the cost of decreased VL accuracy compared to naive fine-tuning.
@@
Table 3: Results for alternatives to naive fine-tuning for transforming Pythia 160M into an MLLM.
(a) Avg. VL performance with varied model size
(b) Linguistic forgetting with varied model size
Figure 3: Vision-language and linguistic forgetting for varying model sizes. We evaluate the average VL performance and linguistic forgetting for Pythia models in the 160M to 2.8B scale, after training on LLaVA. In Figure 3(b), ★ refers to the 0 mark where no linguistic forgetting occurs. Negative forgetting refers to a positive backward transfer in NL performance after multimodal training.
Analyzing the Role of Parameter Count.
In Driess et al. (2023), larger models had less catastrophic forgetting than smaller ones for NLG/NLU, where their 12B parameter model lost $87.3\%$. In comparison, their 562B parameter model only lost $3.9\%$. The opposite result was found in Luo et al. (2023b), where smaller models exhibited less forgetting than larger ones. We analyze this phenomenon in the Pythia family of models, where we evaluate naive fine-tuning and Soft Targets, which was the best method for Pythia (160M). Results are given in Fig. 3. Across scales, Soft Targets has zero or negative linguistic forgetting over the naive fine-tuning method used in the original LLaVA paper, with competitive VL performance compared to naive fine-tuning. Soft Targets achieved positive backward transfer in the low model size regime (0.16-0.41B) and no forgetting for larger sizes (>0.41>0.41B). In contrast, naive fine-tuning had severe forgetting in the 0.16-0.41B regime and reduced forgetting when the model size exceeded 1B parameters.
@@ 4.3 Continually Learning VL Tasks
We next turn to continual learning on each of the VL datasets used to train LLaVA 1.5, where we group the datasets based on the task type. The task order is provided in Table 5, with additional details in Appendix C. Note that there is no evaluation dataset associated with task 2, but we still measure performance on task 1. Given our limited computational budget, we exhaustively studied our CL methods only for the smaller-scale Pythia 410M LLM, and then we evaluated the best-performing mitigation method for all of the Pythia LLMs (from 160M to 1.4B parameters). Our goal in this setting is continually learn new VL tasks, while mitigating both linguistic forgetting as well as forgetting previously learned VL tasks.
@@
Rehearsal $(1\%)$ - 0.58 37.74 10.90 3.47 7.41 3.55 7.65
mSGM + Reh. $(1\%)$ - 1.11 35.28 0.73 12.38 2.25 10.21 2.77
Table 4: Continually learning LLaVA Tasks with Pythia 410M. We report task-wise averaged accuracy and forgetting of each mitigation method across VL and NLU/NLG tasks, where we evaluate all test sets associated with the tasks seen up to the current task.
(a) Task averaged VL Performance with varied model size
(b) Task averaged Linguistic Forgetting with varied model size
Figure 4: Continual Learning LLaVA Tasks: Task averaged Vision-language performance and Linguistic forgetting for varying model sizes. We evaluate mSGM + Rehearsal $(1\%)$ and LLaVA Naive-FT (solid lines for both) on the Continual Learning setup, with varying base LLMs: Pythia models from 160M to 1.4B scale. For each scale, we report the average VL performance or linguistic forgetting across all tasks (task averaged). Performance of joint training (Joint T.) is shown as an upper-bound (dotted lines) for both Naive FT and mSGM + Rehearsal.
@@
Task Type Task Data Size
Pre-Training 1 LLM Pre-Training*
Instruct Tuning 2 $C$-LAION-SBU 558K
LLaVA-Inst, ShareGPT 198K
VQA 3 VQA2 83K
(OE & OK) OKVQA 9K
A-OKVQA 66K
GQA 72K
VQA (OCR) 4 OCRVQA 80K
TextCaps 22K
Referential 5 RefCOCO 48K
Grounding VisualGenome 86K
Table 5: Continual LLaVA Setup. The LLaVA 1.5 data mixture is split into groups of vision-language (VL) tasks. OE & OK refers to open-ended and outside-knowledge VQA.
Results for Pythia 410M are given in Table 4. In terms of NLU/NLG forgetting, all mitigation methods showed efficacy in reducing forgetting across the sequence of tasks. Overall, mSGM with and without rehearsal achieves the least linguistic forgetting while maintaining the highest VL performance across all baselines. Regarding VL performance, naive fine-tuning achieves better performance at the cost of high linguistic forgetting. In contrast, mSGM achieves competitive VL performance and sometimes even surpasses it, e.g., for RefCOCO, an especially challenging VL task. We also report raw accuracies across all tasks in Appendix D.5.
Model scaling results.
Model scaling results are given in Fig. 4. We compare mSGM with Rehearsal, the best method identified for Pythia 410M with naive fine-tuning. Naive fine-tuning leads to a sharp and consistent drop in task averaged NLU/NLG performance, across all model scales. In contrast, mSGM with Rehearsal has little to no NLU/NLG forgetting across all tasks and multiple model scales. For Pythia 1B, mSGM with Rehearsal achieves positive backward transfer for NLU/NLG datasets as the VL tasks are learned, which is rarely observed in CL. For the 160M parameter model, naive fine-tuning suffers large amounts of forgetting compared to mSGM. Across all scales, naive fine-tuning exhibits greater losses of NLU/NLG performance, unlike mSGM. In general, VL performance for mSGM rivals naive fine-tuning across scales. And in the case of the largest 1.4B scale, exceeds the task averaged performance of naive fine-tuning.
@@ 5 Discussion
This work presents one of the first studies of linguistic forgetting in MLLMs, particularly for open-source models with modest parameter counts (<7<7B). We show that the degree of linguistic forgetting typically reduces with model scale, and is far more severe for NLG tasks compared to NLU. Additionally, we demonstrate that instruction-tuned base LLMs are more robust to linguistic forgetting from multimodal training than non instruction-tuned base LLMs. We pioneer treating MLLM creation as a CL problem, and show that CL methods effectively mitigate linguistic forgetting while minimally hindering VL accuracy. In our experiments, our best mitigation method far outperforms the naive LLaVA recipe in terms of linguistic forgetting, while maintaining competitive VL performance. We show that this benefit holds across model scales. We pioneer CL for MLLMs, and establish strong baselines for this task. In fact, we observe that besides maintaining competitive VL performance with naive LLaVA training, our mitigation methods achieve near zero and in some cases even negative linguistic forgetting. This suggest that our mitigation methods achieve a positive backward transfer of linguistic ability after multimodal training. Given the essential nature of multimodal abilities for many applications, our research highlights that naive multimodal fine-tuning can significantly degrade prior linguistic abilities. Our findings underscore the importance of developing more robust and capable mitigation approaches, showcasing the applicability of CL techniques to the fine-tuning of foundation models. We aim to inspire future research in this direction, ultimately contributing to the advancement of more resilient and versatile MLLMs.
6 Future Work
In this work, we rigorously studied catastrophic forgetting in MLLMs trained with the LLaVA recipe. While our findings provide valuable insights, several avenues remain open for future exploration.
We studied LLaVA, but we assume our findings would apply to similar models (Liu et al., 2024; 2023a; Lin et al., 2023a). Of the PET methods, we only studied LoRA; however, other PET techniques may be more effective (Han et al., 2024). Future work could study other methods and investigate their ability to reduce forgetting in MLLMs.
We experimented with 9 MLLMs and our experiments were computationally demanding. Therefore, we could not study all of the methods for every LLM, meaning that using the best approach for only a small LLM may not have ultimately been the best for every LLM. Likewise, our limited computational resources meant we could not study variability in performance for all permutations of the tasks in our CL experiments. Future work might investigate how learning more tasks affects forgetting in MLLMs.
We only studied models that integrate vision and language, but some MLLMs include additional modalities (Driess et al., 2023). It is not clear if our methods will be as effective for those. It would be interesting to study MLLMs that also have multimodal outputs for multimodal dialogue in CL settings. FROMAGe studied this in a non-CL setting with frozen LLM (Koh et al., 2023). Exploration of different CL methods is an exciting future direction to build a multimodal chatbot or agent that can continually and seamlessly interact with users.
Following the LLaVA training recipe, we only trained models with a single pass through the VL datasets. We hypothesize that the larger MLLMs may be under-trained based on our results, considering that their VL performance plateaued compared to smaller models. Adopting this standard does help to ensure a consistent and fair analysis; however, for practical purposes of creating better MLLMs, it would be ideal to maximize VL performance while minimizing NLU/NLG forgetting.
Due to our limited compute budget, we could not study MLLMs with more than 7B parameters. It would be interesting to examine how model size impacts catastrophic forgetting in larger LLMs. For instance, PaLM-E (12B) exhibits 87.387.3% forgetting whereas the largest model, PaLM-E (562B), exhibits only 3.93.9% forgetting. Scaling our methods to larger models is an exciting direction for future work.
@@ Acknowledgments
  • Appendix A includes reproducibility and code release information.
  • Appendix B provides details on compute resources (GPU) and training time.
  • Appendix C provides details on the datasets used in this paper.
@@
  • Appendix D summarizes findings of additional supporting experiments and ablation studies.
Appendix A Reproducibility and Code Release
All datasets and pre-trained models utilized in this work are widely recognized and publicly available. To facilitate replication of our experiments, we will create a GitHub repository with codes and a dedicated website. The code will be released under an MIT license. We will also release our trained MLLMs (weights and configurations). These resources will provide the exact sequence of examples used for each continual learning experiment. Additionally, the Appendix contains a comprehensive description of the hyperparameters and optimizer settings employed in each experiment.
Appendix B Hardware and Computational Budget
We trained a total of 9 MLLMs on 6 datasets, over 170 experiments on 4 A100 40GB GPU, with an aggregated compute time of $\sim 3400$ (wall clock) hours ($\sim 141$ days).
B.1 Hyperparameter Search
B.1.1 LLaVA1.5 Setting
@@ B.1 Hyperparameter Search
To reproduce and directly compare against the LLaVA 1.5 protocol, we keep the explicit training configurations the same (as mentioned in Appendix 3.4).
B.1.2 LoRA
@@
We train several different LoRA settings on the smallest Pythia-160M LLM for tractability and compare the resulting VL and NLU/NLG performances. We vary the a) target modules (between 1) all linear layers, and 2) key, query, and value projection layers only), b) LoRA ranks in the range of 1/41/21/4-1/2 of the full rank of the model, rank stabilized LoRA (Kalajdzievski, 2023), and larger alpha values ($16$ instead of $8$ as default). Table 6 shows the comparisons by training on the LLaVA setting.
@@
B.1.3 Soft Targets
We train the LLaVA recipe with the soft targets with varying alpha $\alpha\in\{0.001,0.01,0.1\}$, and report results in Figure  7. The Pythia 160M model is used for this tuning.
@@ B.1.3 Soft Targets
Appendix C Dataset Preparation
C.1 Continual LLaVA
All datasets are primarily in the English language. However, some small samples of other languages such as Chinese can be found in the ShareGPT subset of LLaVA Instruct (Liu et al., 2023b), for instance. The preprocessing steps required for splitting the LLaVA 1.5 Instruct json into component datasets, to build our continual learning setting, are detailed below: The primary LLaVA instruct json llava_v1_5_mix665k.json (Liu et al., 2023b) is split into component datasets based on the data entries as follows:
  • LLaVA Instruct (Liu et al., 2023b): The first 157,712 entries are directly assigned to the LLaVA subset, per manual verification.
  • ShareGPT (SG) (Liu et al., 2023a): Entries without an associated image are categorized under ShareGPT.
  • GQA (Hudson & Manning, 2019), OCR-VQA, TextVQA (Singh et al., 2019), Visual Genome (VG) (Krishna et al., 2017): Entries associated with images from the GQA, OCR-VQA, TextVQA, and VG_100K are mapped correspondingly to their respective datasets.
  • COCO-Based Datasets: For images related to the COCO dataset, further categorization is performed based on the prompts:
    • OKVQA (Marino et al., 2019): If the prompt contains the phrase “Answer the question using a single word or phrase” and the image ID matches those in the OKVQA set.
    • VQA-V2 (Goyal et al., 2017): If the above condition is not met, the entry is categorized under VQA-V2.
    • A-OKVQA (Schwenk et al., 2022): If the prompt contains multiple choice options (‘A.’, ‘B.’, ‘C.’).
    • RefCOCO (Kazemzadeh et al., 2014): If the prompt requests bounding box coordinates or a short description of a region.
The splitting process categorizes the dataset into distinct subsets as shown in Table 8, which are saved as separate JSON files in the specified directory. The number of entries in each subset is printed for verification.
Table 8: Dataset Split generated based on LLaVA 1.5 Instruct composition Liu et al. (2023a).
Subset Sample Size
LLaVA 157,712
SG 40,688
GQA 72,140
OKVQA 8,994
OCRVQA 80,000
A-OKVQA 66,160
TextCaps 21,953
RefCOCO 48,447
VG 86,417
VQA-V2 82,787
After splitting, the datasets are then combined, based on the grouping in Table 5, to form the respective sequence of Tasks 1 to 5.
C.2 Evaluation Dataset Splits